OpenStudy (anonymous):

70% of a number is 98. What is 90% of that same number? A. 63 B. 68.6 C. 88.2 D. 126

OpenStudy (anonymous):

0.7x=98 0.9x=? 0.7x=98 x=98/0.7=140 0.9*140=126

OpenStudy (anonymous):

88% of what number is 66?

OpenStudy (anonymous):

0.88x=66 solve equation and youll get the answer